# DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures

### v1

@@ 4 Evaluation
In this section, we present experimental results evaluating all nine LLMs and highlight their strengths and weaknesses across structural reasoning tasks on DSR-Bench and DSR-Bench-challenge (Section 4.1), DSR-Bench-spatial (Section 4.2), and DSR-Bench-natural (Section 4.3).
@@ 4.1 Can LLMs understand and manipulate data structures?
In this section, we present our experimental results of LLMs on DSR-Bench (including DSR-Bench-challenge). In Section 4.1.1, we discuss the results of the five instruction-tuned models, and in Section 4.1.2, we highlight key observations on the four reasoning models.
| Relationship | Data Structure | o4- mini | Gemini- 2.5-Pro | Claude-3. 7-Sonnet | DeepSe ek-R1 | GPT- 4.1 | Gemini- 2.0-Flash | Claude-3. 5-Sonnet | DeepSe ek-V3 | Llama 3.3-70B |
| Linear | Array | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 |
| Category avg. | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 | |
| Temporal | Stack | 1.00 | 1.00 | 0.97 | 0.99 | 0.55 | 0.36 | 0.99 | 0.41 | 0.04 |
| Queue | 1.00 | 1.00 | 1.00 | 0.98 | 0.55 | 0.36 | 0.99 | 0.43 | 0.25 | |
| LRU | 1.00 | 1.00 | 0.30 | 0.16 | 0.85 | 0.78 | 0.82 | 0.01 | 0.50 | |
| Priority Queue | 0.55 | 0.51 | 0.28 | 0.65 | 0.25 | 0.16 | 0.33 | 0.20 | 0.08 | |
| Category avg. | 0.89 | 0.84 | 0.64 | 0.69 | 0.55 | 0.42 | 0.79 | 0.26 | 0.22 | |
| Associative | Hashmap | 0.51 | 0.64 | 0.63 | 0.33 | 0.06 | 0.10 | 0.16 | 0.00 | 0.00 |
| Trie | 0.68 | 0.80 | 0.08 | 0.49 | 0.18 | 0.17 | 0.49 | 0.02 | 0.00 | |
| Suffix Tree | 0.73 | 0.85 | 0.91 | 0.96 | 0.00 | 0.01 | 0.08 | 0.67 | 0.00 | |
| Skip List | 0.62 | 0.73 | 0.75 | 0.69 | 0.07 | 0.06 | 0.42 | 0.02 | 0.01 | |
| Category avg. | 0.63 | 0.76 | 0.59 | 0.62 | 0.08 | 0.09 | 0.29 | 0.18 | 0.00 | |
| Hierarchical | BST | 0.86 | 0.86 | 0.64 | 0.73 | 0.59 | 0.43 | 0.71 | 0.58 | 0.34 |
| Heap | 0.68 | 0.61 | 0.40 | 0.48 | 0.20 | 0.10 | 0.27 | 0.15 | 0.07 | |
| RB tree | 0.65 | 0.77 | 0.30 | 0.62 | 0.12 | 0.12 | 0.46 | 0.09 | 0.05 | |
| B+ tree | 0.97 | 0.95 | 0.38 | 0.31 | 0.23 | 0.12 | 0.23 | 0.08 | 0.01 | |
| K-D Tree | 0.47 | 0.71 | 0.00 | 0.45 | 0.00 | 0.01 | 0.03 | 0.00 | 0.00 | |
| K-D Heap | 0.10 | 0.05 | 0.05 | 0.11 | 0.04 | 0.04 | 0.05 | 0.03 | 0.01 | |
| Category avg. | 0.64 | 0.78 | 0.33 | 0.45 | 0.20 | 0.14 | 0.34 | 0.16 | 0.08 | |
| Network | Graph | 0.87 | 0.94 | 0.11 | 0.67 | 0.15 | 0.05 | 0.06 | 0.06 | 0.02 |
| DSU | 0.98 | 1.00 | 0.99 | 1.00 | 0.02 | 0.00 | 0.02 | 0.82 | 0.00 | |
| Geom Graph | 0.16 | 0.13 | 0.02 | 0.13 | 0.07 | 0.02 | 0.02 | 0.01 | 0.02 | |
| Category avg. | 0.67 | 0.69 | 0.38 | 0.60 | 0.08 | 0.03 | 0.03 | 0.30 | 0.01 | |
| Hybrid | Bloom Filter | 0.61 | 0.89 | 0.16 | 0.74 | 0.04 | 0.04 | 0.04 | 0.04 | 0.02 |
| DAWG | 0.25 | 0.07 | 0.06 | 0.14 | 0.05 | 0.06 | 0.07 | 0.06 | 0.01 | |
| Category avg. | 0.43 | 0.48 | 0.11 | 0.44 | 0.05 | 0.05 | 0.06 | 0.05 | 0.02 | |
| Score | Overall avg. | 0.72 | 0.76 | 0.51 | 0.63 | 0.31 | 0.27 | 0.41 | 0.32 | 0.17 |
@@ 4.1.1 Instruction-tuned models
As shown in Table 1, these models show sharp accuracy drops in tasks involve elements with multiple attributes. For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority. Similarly, in the hashmap task, manual inspection of errors shows that models confuse keys and values, delete the wrong items, or hallucinate entries. These results reveal a key limitation for real-world deployment, where managing entities with multiple interacting properties, such as deadlines, priorities, or key-value records, is common.
@@
We see from Table 1 that, while models perform reliably on binary search trees, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across multiple ancestral levels. Performance declines further on B+ trees, which require reasoning over wider spans of child pointers, and on graph traversal tasks involving many-to-many relationships. These results underscore a persistent limitation: instruction-tuned LLMs struggle to maintain and manipulate context beyond local neighborhoods.
@@
As shown in Table 2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial. Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures. A representative case is suffix-tree construction: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section A.5).
@@ 4.1.2 Reasoning models
As shown in Table 1, reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures. However, overall score remains below 0.5 on DSR-Bench-challenge (Table 3), in particular for long tasks and complex data structures. For example, the highest score on heap is only 0.62, despite it being a widely used structure in scheduling and resource allocation. Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Appendix A.4). To further probe these limitations, we introduce DSR-Bench-spatial, described in Section 4.2.
@@
We observe that while GPT-4.1 scores 0.66 on the binary search tree depth task (short), o4-mini performs much worse, with a score of just 0.01 (Section A.4). Manual inspection of incorrect outputs reveals that o4-mini consistently interprets depth as the number of edges rather than nodes, despite explicit instructions. Similarly, in an ablation on K-D Heap (Table 5), switching from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons. Directly quering o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys. These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors, limiting their reliability on tasks with strict user-defined constraints.
@@
4.2 Can LLMs reason structurally on spatial data?
Real-world data is often represented in multi-dimensional feature spaces. To assess whether LLMs can reason over such spatial data, we extend the benchmark with DSR-Bench-spatial, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space. These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics. Given task complexity, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps for Section 4.2 and Section 4.3.
@@ 4.2 Can LLMs reason structurally on spatial data?
As shown in Table 5, accuracy consistently drops for both models as dimensionality increases. Higher-dimensional data challenges models to reason over more complex distance metrics and partitions, limiting their effectiveness in spatial tasks. For instance, K-D trees are commonly used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision [40]. However, a notable exception is the Geometric Graph, where 2D data performs best, likely because it is more commonly found in textbook examples used as training data.
@@
We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data. Specifically, we test K-D tree construction tasks using three non-uniform distributions from scikit-learn [37]: circles, moons, and blobs (examples shown in Figure 3, Section A.6). As shown in Table 6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in training data. Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning. In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts.
4.3 Can LLMs reason structurally on natural language?
While the previous sections evaluated LLMs on canonical data structures, real-world use-cases are often expressed in natural language. To bridge this gap, we extend the benchmark with DSR-Bench-natural, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
@@ 4.3 Can LLMs reason structurally on natural language?
We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling). Synthetic data follows the same distributions as Section 2.2, with realistic substitutions (e.g., names for integers). Each scenario was written by humans, paraphrased by GPT-4o. All prompts were reviewed by three annotators for clarity and unambiguity. Details of prompt generation can be found in Section A.7.
@@
As shown in Table 7, performance drops when tasks are described in natural language from formal descriptions, despite identical operation distributions. This suggests that even reasoning models struggle to apply their abilities in language-rich, real-world contexts. Bridging this gap is crucial for reliable deployment and presents a key direction for future research. Furthermore, the higher accuracy on formal descriptors may stem from LLMs’ exposure to textbook-style patterns, where integers and explicit syntax are common.
5 Related Works
LLM reasoning benchmarks
Many benchmarks have been proposed to evaluate the capabilities and limitations of LLMs (see Section A.1 for a detailed review). Reasoning-focused benchmarks have primarily been domain-specific, targeting areas such as math [5, 28, 25], coding challenges [50, 16, 3, 58], STEM question answering [15], and logic puzzles [50, 11]. These often require long, complex responses involving intertwined reasoning steps, and rely on human or LLM-based judgment for evaluation [4, 10, 54]. In contrast, we take a more foundational approach: testing whether LLMs can understand and manipulate data relationships, which is a core capability that underlies reasoning in math and programming, isolating structural reasoning from other complexities.
Algorithmic reasoning with LLMs
Prior work on algorithmic reasoning has primarily focused on arithmetic tasks and length generalization in transformers [59, 60, 21], or on specific graph problems such as cycle detection and connectivity [48, 8]. These efforts offer limited coverage of broader data structure reasoning. One closely related work, CLRS-Text [32], evaluates whether LLMs can simulate 30 algorithms from the CLRS benchmark suite [47, 6] via textual inputs. Rather than testing whether LLMs can replicate specific algorithms, we evaluate their ability to reason over data structures, which are core abstractions and prerequisites for algorithmic execution. This offers a more general and interpretable assessment of algorithmic reasoning, revealing not only whether models succeed, but also which types of structural reasoning they lack.
6 Conclusion, limitations, and future work
We introduce DSR-Bench, the first benchmark for evaluating LLMs’ structural reasoning through canonical data structures. DSR-Bench enables fine-grained, controlled, and objective evaluation using synthetic data and an automated scoring pipeline. Benchmarking nine state-of-the-art LLMs reveals key limitations across both instruction-tuned and reasoning models. Notably, DSR-Bench-challenge remains difficult even for the strongest models. In addition, we propose two extensions, DSR-Bench-spatial and DSR-Bench-natural, to evaluate spatial reasoning and natural language understanding of data structures. These results uncover further gaps, suggesting current LLMs are not yet ready for structure-dependent, real-world tasks.
Limitation and future work
While our scoring system compares the final state of the data structure, additional insights could be gained by evaluating models’ intermediate thought traces. However, such traces are unavailable from some providers and can be costly to retrieve. As interfaces grow more transparent, we hope to incorporate such signals to enrich error analysis and better understand model reasoning. Additionally, while DSR-Bench covers a wide range of structures, expanding it to cover emerging representations, such as graph-probabilistic hybrid structures, would further broaden its scope. Our benchmark is modular and extensible, and we welcome community contributions to advance these directions and help develop an even more comprehensive evaluation framework.
Acknowledgments and Disclosure of Funding

### v2

@@ 4 Evaluation
Models
We evaluate ten state-of-the-art LLMs across two categories: instruction-tuned models and reasoning models. We select the latest flagship models from each major provider, with each problem evaluated three times (166,230 total evaluations). Instruction-tuned models are widely deployed due to their efficiency and scalability, making their structural reasoning capabilities practically important: we cover GPT-4.1-2025-04-14 (OpenAI et al., 2024), Llama3.3-70B (Grattafiori et al., 2024), Gemini-2.0-Flash-001 (Team et al., 2024), Claude-3-5-Sonnet-20241022 (Anthropic, 2024), and DeepSeek-V3 (DeepSeek-AI et al., 2025). Reasoning models are explicitly trained for complex, multi-step reasoning: we evaluate GPT-5-2025-08-07 with medium thinking effort (OpenAI, 2025), o4-mini-2025-04-16 (OpenAI, 2024), Gemini-2.5-Pro (stable) (Team et al., 2024), Claude-3-7-Sonnet-20250219 (Anthropic, 2024), and DeepSeek-R1 (DeepSeek-AI et al., 2025).
Prompting strategies
We also study the impact of prompting strategies on data structure tasks. Unlike reasoning models with internal multi-step inference (e.g., reasoning tokens), instruction-tuned models are particularly sensitive to prompt formulation. We evaluate five prompting strategies: (i) Stepwise, which adds a “steps” field to the output JSON schema; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single example with intermediate reasoning steps; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text. See Section˜A.3 for examples.
@@ 4.1 Can LLMs understand and manipulate data structures?
In this section, we present our main experimental results on DSR-Bench, including main and its challenging subset challenge. We discuss insights from five instruction-tuned models in Section˜4.1.1 and from five reasoning models in Section˜4.1.2.
| Relationship | Data Structure | GPT-5 (med) | o4- mini | Gemini- 2.5-Pro | Claude-3. 7-Sonnet | DeepSe ek-R1 | GPT- 4.1 | Gemini- 2.0-Flash | Claude-3. 5-Sonnet | DeepSe ek-V3 | Llama 3.3-70B |
| Linear | Array | 1.00 | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 |
| Category avg. | 1.00 | 1.00 | 1.00 | 0.99 | 0.99 | 0.94 | 0.90 | 0.96 | 0.98 | 0.69 | |
| Temporal | Stack | 1.00 | 1.00 | 1.00 | 0.97 | 0.99 | 0.55 | 0.36 | 0.99 | 0.41 | 0.04 |
| Queue | 1.00 | 1.00 | 1.00 | 1.00 | 0.98 | 0.55 | 0.36 | 0.99 | 0.43 | 0.25 | |
| LRU | 1.00 | 1.00 | 1.00 | 0.30 | 0.16 | 0.85 | 0.78 | 0.82 | 0.01 | 0.50 | |
| Priority Queue | 0.52 | 0.55 | 0.51 | 0.28 | 0.65 | 0.25 | 0.16 | 0.33 | 0.20 | 0.08 | |
| Category avg. | 0.88 | 0.89 | 0.84 | 0.64 | 0.69 | 0.55 | 0.42 | 0.79 | 0.26 | 0.22 | |
| Associative | Hashmap | 0.87 | 0.51 | 0.28 | 0.63 | 0.33 | 0.06 | 0.10 | 0.16 | 0.00 | 0.00 |
| Trie | 0.94 | 0.68 | 0.62 | 0.08 | 0.49 | 0.18 | 0.17 | 0.49 | 0.02 | 0.00 | |
| Suffix Tree | 0.98 | 0.73 | 0.90 | 0.91 | 0.96 | 0.00 | 0.01 | 0.08 | 0.67 | 0.00 | |
| Skip List | 0.68 | 0.62 | 0.78 | 0.75 | 0.69 | 0.07 | 0.06 | 0.42 | 0.02 | 0.01 | |
| Category avg. | 0.87 | 0.63 | 0.65 | 0.59 | 0.62 | 0.08 | 0.09 | 0.29 | 0.18 | 0.00 | |
| Hierarchical | BST | 1.00 | 0.86 | 0.97 | 0.64 | 0.73 | 0.59 | 0.43 | 0.71 | 0.58 | 0.34 |
| Heap | 0.61 | 0.68 | 0.38 | 0.40 | 0.48 | 0.20 | 0.10 | 0.27 | 0.15 | 0.07 | |
| RB tree | 0.76 | 0.65 | 0.49 | 0.30 | 0.62 | 0.12 | 0.12 | 0.46 | 0.09 | 0.05 | |
| B+ tree | 0.98 | 0.97 | 0.97 | 0.38 | 0.31 | 0.23 | 0.12 | 0.23 | 0.08 | 0.01 | |
| K-D Tree | 0.85 | 0.47 | 0.59 | 0.00 | 0.45 | 0.00 | 0.01 | 0.03 | 0.00 | 0.00 | |
| K-D Heap | 0.10 | 0.10 | 0.10 | 0.05 | 0.11 | 0.04 | 0.04 | 0.05 | 0.03 | 0.01 | |
| Category avg. | 0.72 | 0.64 | 0.58 | 0.33 | 0.45 | 0.20 | 0.14 | 0.34 | 0.16 | 0.08 | |
| Network | Graph | 0.96 | 0.87 | 0.78 | 0.11 | 0.67 | 0.15 | 0.05 | 0.06 | 0.06 | 0.02 |
| DSU | 1.00 | 0.98 | 0.97 | 0.99 | 1.00 | 0.02 | 0.00 | 0.02 | 0.82 | 0.00 | |
| Geom Graph | 0.16 | 0.16 | 0.13 | 0.02 | 0.13 | 0.07 | 0.02 | 0.02 | 0.01 | 0.02 | |
| Category avg. | 0.71 | 0.67 | 0.63 | 0.38 | 0.60 | 0.08 | 0.03 | 0.03 | 0.30 | 0.01 | |
| Hybrid | Bloom Filter | 0.77 | 0.61 | 0.86 | 0.16 | 0.74 | 0.04 | 0.04 | 0.04 | 0.04 | 0.02 |
| DAWG | 0.34 | 0.25 | 0.23 | 0.06 | 0.14 | 0.05 | 0.06 | 0.07 | 0.06 | 0.01 | |
| Category avg. | 0.56 | 0.43 | 0.55 | 0.11 | 0.44 | 0.05 | 0.05 | 0.06 | 0.05 | 0.02 | |
| Score | Overall avg. | 0.79 | 0.72 | 0.71 | 0.51 | 0.63 | 0.31 | 0.27 | 0.41 | 0.32 | 0.17 |
@@ 4.1.1 Instruction-tuned models
As shown in Table˜1, these models show sharp accuracy drops in tasks involving elements with multiple attributes. For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority. Similarly, in the hashmap task, manual inspection of errors shows that models confuse keys and values, delete the wrong items, or hallucinate entries. These results reveal a key limitation for real-world deployment, where managing entities with multiple interacting properties, such as deadlines, priorities, or key-value records, is common.
@@
We see from Table˜1 that, while models perform reliably on BST, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across ancestral levels. Performance declines further on B+ trees, which requires reasoning over wider spans of child pointers, and on graph traversal tasks with many-to-many relationships. Manual inspection of reasoning traces reveals failures to retain earlier information: in Geometric Graph-long, all GPT-4.1 errors stemmed from dropping nodes during intermediate steps.
@@
As shown in Table˜2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial. Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section˜A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures, but require careful design. A representative case is suffix-tree: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section˜A.5). We include more analysis of CoT to provide meaningful takeaways for practitioners in Section˜A.6.
@@ 4.1.2 Reasoning models
From Table˜1, we see reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures. However, the overall score remains below 0.5 on challenge in Table˜3, in particular for long tasks and complex data structures. For example, the highest score on skip list is only 0.61, despite its prevalence in introductory-level textbooks and its wide use in dictionaries and maps. Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Section˜A.4). To further probe these limitations, we introduce spatial, described in Section˜4.2.
@@
In an ablation on K-D Heap in Table˜5, switching the tie-breaking rule from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons. Directly querying o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys. These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors learned from training, limiting their reliability on tasks with user-defined constraints.
4.2 Can LLMs reason structurally on spatial data?
Real-world data is often represented in high-dimensional feature spaces. To assess whether LLMs can reason over such spatial data, we extend the benchmark with the spatial suite, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space. These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics. Given the complexity of these tasks, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps.
@@ 4.2 Can LLMs reason structurally on spatial data?
As shown in Table˜5, accuracy declines for both models as dimensionality increases. Higher-dimensional data challenges models with more complex computation over distance metrics and partitions, limiting their effectiveness in spatial tasks. For instance, K-D trees are widely used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision (Silpa-Anan & Hartley, 2008). Interestingly, 2D outperforms 1D in Geometric Graph, likely due to its more common presence in training data from textbooks.
@@
We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data. We test K-D tree construction tasks using three non-uniform distributions from scikit-learn (Pedregosa et al., 2011): circles, moons, and blobs (illustration in Figure˜3, Section˜A.7). As shown in Table˜6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in the training data. Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning. In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts. A more in-depth inspection of errors and discussion on root causes can be found in Section˜A.7.
4.3 Can LLMs reason structurally on natural language tasks?
While the previous sections evaluated LLMs on canonical data structures, real-world use cases are often expressed in natural language. To bridge this gap, we extend the benchmark with the natural suite, which embeds data structure tasks in narrative, real-world contexts, allowing us to test whether LLMs generalize structural reasoning beyond formal descriptions.
@@ 4.3 Can LLMs reason structurally on natural language tasks?
We design three real-world scenarios that implicitly require data structures: queue (children buying ice cream), binary search tree (clinic appointments), and graph (galaxy traveling game). Synthetic data follows the same distributions as Section˜3.3, with realistic substitutions (e.g., names for integers). Each scenario was written by humans and paraphrased by GPT-4o. All prompts were reviewed by three annotators for clarity and unambiguity. Details are supplemented in Section˜A.8.
@@
As shown in Table˜7, performance drops when tasks are described in natural language compared to formal descriptions, despite identical problem distribution. The higher accuracy on formal descriptors may stem from training on textbook-style patterns, where integers and explicit syntax are common. This observation suggests that even reasoning models struggle to apply reasoning in language-rich, real-world contexts. Bridging this gap is crucial for reliable deployment and presents a key direction for future research.
4.4 Can LLMs reason structurally with code generation?
As motivated in Section˜1 and Section˜2, our benchmark targets general reasoning independent of code execution or tool use. Nonetheless, to assess whether code generation provides any benefit, we run ablations on six models using the code suite across three modes: (i) CodeOnly, where models generate Python code executed by an external interpreter; (ii) CodeEnforce, where models must write code and reason through its execution internally without relying on an interpreter; and (iii) CodeMaybe, similar to CodeEnforce but makes code generation optional. Full details on the experimental setup and results are provided in Section˜A.9.
| Mode | GPT- 4.1 | o4- mini | Gemini- 2.0-Flash | Gemini- 2.5-Pro | Claude-3. 5-Sonnet | Claude-3. 7-Sonnet |
| Default | 0.40 | 0.73 | 0.38 | 0.55 | 0.44 | 0.55 |
| CodeMaybe | 0.38 | 0.76 | 0.41 | 0.55 | 0.43 | 0.57 |
| CodeEnforce | 0.38 | 0.75 | 0.41 | 0.53 | 0.42 | 0.55 |
| CodeOnly | 0.95 | 0.82 | 0.44 | 0.57 | 0.74 | 0.87 |
| o4- mini | Gemini- 2.5-pro | Claude-3. 7-Sonnet | |
| Geom Graph | 0.99 | 0.98 | 0.96 |
| DAWG | 0.56 | 0.90 | 0.89 |
| Graph-Natural | 0.69 | 0.56 | 0.23 |
Models cannot reason over generated code.
In Table˜9, performance in CodeMaybe and CodeEnforce matches the default setup, showing that writing code offers little benefit over natural language reasoning when models must internally simulate execution. This reinforces our central claim: LLMs remain limited in their ability to perform structural reasoning, even when guided by their own code.
Code helps only with standard tasks and fails on natural language ones.
As shown in Table˜9, with an external interpreter in CodeOnly, models perform well on Geom Graph, a standard structure in computer graphics with widely available implementations. In contrast, they struggle on the less familiar DAWG, where custom constraints enforce unambiguous outputs, suggesting reliance on memorized solutions rather than genuine reasoning. Performance drops further on Graph-Natural, where models default to brittle pattern matching (e.g., rigidly mapping “A space tunnel links A and B” to “G.add_edge(A, B)”) but failing to understand paraphrases like “Couriers frequently travel the tunnel connecting A to B.” These results highlight the fragility of structural reasoning under natural language ambiguity, even with code generation and external execution.
5 Discussion and conclusion
Can LLMs reason structurally? Through DSR-Bench, we provide a systematic answer: not yet. Instruction-tuned models struggle with multi-attribute reasoning (e.g., database indexing) and multi-hop reasoning (e.g., trip planning), while reasoning models achieve only 0.498 accuracy on complex structures and can ignore user-defined constraints. These limitations highlight the need for architectures that support precise function computation, memory mechanisms, and the flexibility to adapt to personalized requests. Evaluations on high-dimensional data (spatial) and natural language scenarios (natural) further reveal gaps between current reasoning capabilities and real-world readiness. Code generation modes (code) show that models cannot reliably reason over their own code, often reverting to memorized patterns or brittle mappings even with external execution.
DSR-Bench provides a systematic framework for evaluating algorithmic reasoning through the lens of structural reasoning. It provides the community with a powerful diagnostic tool: researchers can pinpoint failure modes, test targeted improvements, and measure progress on specific relationship types, paving the way for algorithm-centric model design as an alternative to scaling alone. It also raises new questions: Can LLMs dynamically choose reasoning strategies? Where and why do errors arise in intermediate steps? We invite the community to use DSR-Bench to explore these directions.
Acknowledgments